A desuperheater can be installed on your refrigeration system to obtain free hot water. You estimate you can heat 150,000 liters of water per year from 15°C (city water entering temperature) to 60°C. This will replace a natural gas water heater that operates at 80% efficiency. You pay $10.00 per GJ for your gas. The unit has an economic life of 5 years. How much can you pay for that unit and obtain a return of 12%?
Answer #1

Hear energy required to heat 150,000 liters of water per year from 15°C (city water entering temperature) to 60°C = mass x specific heat x temperature difference = 150,000 x 4186 x (60 - 15) =  28,255,500,000 Joules =  28.2555 GJ

The natural gas water heater is operating at 80% efficiency.

Hence, energy consumed by the natural gas water heater =  28.2555 / efficiency =  28.2555 / 80% =  35.3194 GJ

Annual electricity bill saved = annuity cash inflow = C = 10 x  35.3194 = $  353.1938

N= economic life = 5 years

Return, R = 12%

Hence, maximum we can pay for the new unit = PV of all the electricity bills saved over next five years

= C / R x [1 - (1 + R)-N] =  353.1938 /12% x [1 - (1 + 12%)-5] = $  1,273.1844